# “Quantum Geometric Nesting” and Solvable Model Flat-Band Systems

### v1

@@ Abstract
The properties of a correlated electronic system are largely determined by its electronic structure. One such familiar manifestation is the role of Fermi surface nesting conditions: if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e. $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order. If the overlap occurs under a momentum-space inversion about a certain wavevector $\bm{Q}/2$ (i.e. $\bm{k}\to\bm{Q}-\bm{k}$), the favored order is superconductivity (SC) with particle-particle (p-p) fermion bilinear order and a Cooper pair center of mass momentum $\bm{Q}$. (Such perfect nesting always exists at $\bm{Q}=0$ in time-reversal symmetric systems, where it leads to the familiar Cooper instability.) The thinking behind Fermi surface nesting is based on weak coupling mean-field analysis, in which ordering tendencies are diagnosed by divergent bare susceptibilities in corresponding channels. These concepts are indeed convenient and powerful (see e.g. Refs. Whangbo et al. (1991); Chuang et al. (2001); Kivelson et al. (2003); Shen et al. (2005); Lee et al. (2006); Johannes and Mazin (2008); Terashima et al. (2009); Ran et al. (2009); Nandkishore et al. (2012); Feng et al. (2013); Lee (2014); Jin et al. (2022) for typical applications). However, their validity is limited to weakly coupled systems, and they completely neglect another important piece of information in the electronic structure - the quantum geometry Resta (2011) and topology Bradlyn et al. (2017); Po et al. (2017); Kruthoff et al. (2017) encoded in the single-particle wavefunctions. An entirely different concept of “nesting” is thus needed for flat-band systems, whose bandwidths are small compared to the interaction strengths but which often have non-trivial quantum geometry and rich ordering phenomena Peotta and Törmä (2015); Tovmasyan et al. (2016); Kang and Vafek (2019); Rhim and Yang (2019); Hofmann et al. (2020); Bultinck et al. (2020); Kwan et al. (2021a); Khalaf et al. (2021); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Kwan et al. (2021b); Vafek and Kang (2021); Peri et al. (2021); Liu et al. (2021); Rossi (2021); Angeli and MacDonald (2021); Wang et al. (2022); Hu et al. (2022); Huhtinen et al. (2022); Törmä et al. (2022); Chan et al. (2022a); Herzog-Arbeitman et al. (2022a); Mera and Mitscherling (2022); Pan et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Lin et al. (2022); Herzog-Arbeitman et al. (2022a); Peotta (2022); Khalaf et al. (2022); Chan et al. (2022b); Kang and Vafek (2019); Hofmann et al. (2023); Wan et al. (2023); Swaminathan et al. (2023); Hu et al. (2023); Löthman and Black-Schaffer (2017); Chen and Law (2023); Iskin (2023a); Peotta et al. (2023); Iskin (2023b); Tam and Peotta (2023); Zhang et al. (2023); Pan et al. (2023); Crépel et al. (2023); Mao et al. (2023); Kwan et al. (2023); Sahay et al. (2023); Jiang and Barlas (2023); Scheer and Lian (2023a, b); Devakul et al. (2023); Herzog-Arbeitman et al. (2023); Jia et al. (2023); Wu et al. (2023); Chen and Huang (2023).
@@
In this paper, we propose such a concept, which we call “Quantum Geometric Nesting” (QGN), that focuses on the geometric aspects of the electronic structure in flat-band systems. In contrast to Fermi surface nesting, here we start from a strong-coupling perspective in which the band dispersion is assumed negligible in comparison with the interaction strength, and then ask: for a given quantum geometry of the flat bands, are there any natural forms of order? If so, what sort of interactions favor them? We will show that these questions can be addressed by examining whether a set of QGN conditions is satisfied. For “perfectly nested quantum geometry,” we construct ideal Hamiltonians containing arbitrary combinations of an infinite class of engineered short-range two-body interactions that have solvable ground states with fermion bilinear order of a character determined by the quantum geometry. For these ideal Hamiltonians, the few-body excitations are solvable as well, including single-particle (quasi-particle), particle-hole (exciton), and two-particle (Cooper pair) excitations. For the p-p case, we further show that these ideal Hamiltonians host a (fine-tuned) pseudospin $SU(2)$ symmetry, whose stiffness can be inferred from the excitation spectra. These solvable interacting Hamiltonians can either directly describe realistic systems or serve as starting points for various perturbative studies of additional terms. This type of intrinsically strong coupling viewpoint has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al. (2016); Kwan et al. (2021b); Vafek and Kang (2021); Kang and Vafek (2019); Khalaf et al. (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Pan et al. (2022); Lin et al. (2022); Herzog-Arbeitman et al. (2022a); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Crépel et al. (2023); Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023); Kwan et al. (2023), and it is remarkable that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems. More possibilities also exist within the framework of QGN, which are exemplified with an engineered model for pair-density-wave. Extending the construction scheme, we further obtain an SSH-type model with solvable $p$-wave pairing groundstates.
@@
Figure 1: Schematic flowchart illustrating the structure of the theory: Given any electronic structure containing isolated flat bands, one can test whether a QGN condition is satisfied; if it is, one or more fermion bilinear order parameters (OP) will be suggested (Sec. I). For each OP, infinitely many “ideal” interacting Hamiltonians can be systematically constructed (Sec. II), which, after projecting onto the flat-band Hilbert space $\mathcal{H}_{\text{flat}}$, feature solvable ground states with the corresponding order (Sec. III) and solvable aspects of the collective excitation spectrum (Sec. IV). Some examples of electronic structures with perfect QGN can be found in Sec. V. Possible extensions of this approach are commented on in Sec. VI.
@@ I.1 Preliminaries
We consider an electronic structure encoded in a tight-binding model with charge conservation and translation symmetry among $\mathsf{V}$ unit cells in arbitrary dimension:
@@
where $\bm{R}$ indicates the unit cell position, Greek letters in lowercase $\mu,\nu,\dots$ denote a “flavor” index which, for simplicity, includes both orbital and spin indices unless otherwise specified. The annihilation operators for Bloch states with a given flavor index are expressed as $\hat{c}_{\bm{k},\mu}\equiv\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{R}}\hat{c}_{\bm{R},\mu}/\sqrt{\mathsf{V}}$ (periodic gauge) while the single particle Hamiltonian is diagonalized in a band basis,
@@
We will quantify the degree of QGN for flat-band systems by utilizing the geometric information in $U_{\mu n}(\bm{k})$. To do so, we consider an idealized setup where there are $\mathsf{N}_{\text{flat}}$ nearly degenerate and flat bands isolated from all the other bands by a large gap $\Delta$, such that the relevant interaction strength $V$ is small compared to $\Delta$ but large compared to the flat bands’ bare bandwidth $W$,
@@ I.2 General Definitions
In terms of the gauge-invariant matrices $P$ and $Q$, the key information of QGN at a given wavevector $\bm{Q}$ for p-p or p-h channel is encoded in the linear operator
@@
Viewed as a matrix (with $\mu\nu$ a grouped index and $\mu^{\prime}\nu^{\prime}$ the other), $\Pi$ is hermitian and positive semi-definite and can be diagonalized with non-negative eigenvalues.
By definition, an electronic structure has “perfect QGN” at a wavevector $\bm{Q}$ in the p-p or p-h channel when the corresponding $\Pi$ has eigenvalue(s) equal to 00. In other words, a perfect QGN is satisfied if there exist a $\mathsf{N}\times\mathsf{N}$ matrix $N^{\bm{Q}}_{\mu\nu}$, which we call the “nesting matrix”, satisfying (for any $\mu^{\prime}\nu^{\prime}$)
@@
where $F$ is the form factor
@@
One important property (which can serve as an equivalent definition) of QGN is that for arbitrary $\bm{k}$, the form factor $F_{nm}(\bm{k})$ defined by a nesting matrix is a block-diagonal matrix that disconnects flat band indices from the others, i.e. $F_{nm}(\bm{k})=0$ as long as $n,m$ are not simultaneously $\leq\mathsf{N}_{\text{flat}}$ or $>\mathsf{N}_{\text{flat}}$.
For QGN to be non-trivial in the p-p channel, it is additionally required that $N^{\bm{Q}}$ cannot be symmetric since $\hat{O}^{\text{p-p},\bm{Q}}$ vanishes in this case. (It is proven by explicit construction that valid, anti-symmetric solutions can indeed exist for $\Pi$.) On the other hand, the identity matrix is always an allowed solution for Eq. 7 in the p-h case for $\bm{Q}=\bm{0}$, but this is a trivial one in the sense that the resulting $\hat{O}^{\text{p-h},\bm{Q}=\bm{0}}$ is simply the density operator not associated with any broken symmetry. Excluding all the unwanted solutions, for a generic flat-band system, the value of the smallest eigenvalue of $\Pi^{\bm{Q}}$, $\varpi_{0}^{\bm{Q}}$, is a natural dimensionless measure of “nestability” that quantifies the closeness of a system to perfect QGN in the corresponding channel at $\bm{Q}$  [^2] We note that these concepts are spiritually similar to the “ideal band condition” Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) which are concerned with whether a Chern band is suitable for hosting fractional quantum Hall states.. Later in Sec. V we will give several examples of electronic structures with perfect QGN, i.e. $\varpi_{0}^{\bm{Q}}=0$.
@@ II Ideal interacting Hamiltonians
For generic flat-band systems, no set of local hermitian operators that satisfy Eq. 14 exists. However, when $N$ and $U$ in Eq. 8/9 derive from a band geometry with perfect QGN, there is an infinite class of solutions that take a special flavor-momentum separable form
$\hat{S}^{(I)}_{\bm{R}}=\sum_{\bm{R}_{1}\bm{R}_{2};\mu\nu}s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}B^{(I)}_{\mu\nu}\hat{c}^{\dagger}_{\bm{R}+\bm{R}_{1},\mu}\hat{c}_{\bm{R}+\bm{R}_{2},\nu},$ (15)
where $B^{(I)}$ is a hermitian matrix and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=[s^{(I)}_{\bm{R}_{2},\bm{R}_{1}}]^{\star}$ is the spatial coeffient. In order for Eq. 14 to hold, the constraints on the coefficient $s^{(I)}$ and matrix $B^{(I)}$ are rather loose, in that they only need to obey
$s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=\left[s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}\right]^{\star}\mathrm{e}^{-\mathrm{i}(\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}},$ (16)
@@
for the p-h case. We derive in Sec. II B&C of SM a systematic way of constructing all possible $B^{(I)}$ and prove that there are at least $3\mathsf{N}_{\text{flat}}$ of them in the p-p case, and $\mathsf{N}_{\text{flat}}$ of them in the p-h case when $2\bm{Q}=\bm{0}$. (Some explicit worked examples are presented in Sec. V.) In particular, the identity matrix is always a solution for $B$ in the p-h case, which makes $\hat{S}_{\bm{R}}$ a local density operator for the unit cell at $\bm{R}$. Clearly, even given all the constraints, we are still left with an infinite class of possible interaction terms with considerable leeway, at least in choosing $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}$, which thus could substantially overlap with realistic interactions.
One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”. For the special case of Hubbard interactions, a property called the “uniform pairing condition” Tovmasyan et al. (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al. (2022a). It is possible to generalize this condition, which we will exemplify in Sec. V.3. However, these quadratic terms are not of central focus since the current work is concerned with the existence of ideal interacting Hamiltonians for a given set of isolated flat bands, whose quantum geometry is encoded in the projection matrix and is unaffected by these terms to the leading order in $V/\Delta$. An operational workaround is to simply regard these additional quadratic terms as parts of $\hat{H}_{0}$. For more detailed discussions about this nomenclature subtlety, see Sec. II D of SM.

### v2

@@ Abstract
The properties of a correlated electronic system are largely determined by its electronic structure. One such familiar manifestation is the role of Fermi surface nesting conditions Lomer (1962); Fedders and Martin (1966): if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e. $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order. If the overlap occurs under a momentum-space inversion about a certain wavevector $\bm{Q}/2$ (i.e. $\bm{k}\to\bm{Q}-\bm{k}$), the favored order is superconductivity (SC) with particle-particle (p-p) fermion bilinear order and a Cooper pair center of mass momentum $\bm{Q}$. (Such perfect nesting always exists at $\bm{Q}=0$ in time-reversal symmetric systems, where it leads to the familiar Cooper instability.) The thinking behind Fermi surface nesting is based on weak coupling mean-field analysis, in which ordering tendencies are diagnosed by divergent bare susceptibilities in corresponding channels. These concepts are indeed convenient and powerful (see e.g. Refs. Whangbo et al. (1991); Chuang et al. (2001); Kivelson et al. (2003); Shen et al. (2005); Lee et al. (2006); Johannes and Mazin (2008); Terashima et al. (2009); Ran et al. (2009); Nandkishore et al. (2012); Feng et al. (2013); Lee (2014); Jin et al. (2022) for typical applications). However, their validity is limited to weakly coupled systems in which the interaction strength is small compared to the bandwidth. In general, all forms of non-interacting susceptibilities grow as decreasing $T$ with the same asymptotic behavior $\chi\sim 1/T$ for $T$ greater than the bandwidth, which thus indicates comparable ordering tendencies for all possible order parameters. Significantly, this concept, defined solely by the electronic dispersion, overlooks another important piece of information in the electronic structure - the quantum geometry Provost and Vallee (1980); Resta (2011) and topology Bradlyn et al. (2017); Po et al. (2017); Kruthoff et al. (2017) encoded in the single-particle wavefunctions. An entirely different concept of “nesting” is thus needed for flat-band systems, whose bandwidths are small compared to the interaction strengths but which often have non-trivial quantum geometry and rich ordering phenomena Peotta and Törmä (2015); Tovmasyan et al. (2016); Kang and Vafek (2019); Rhim and Yang (2019); Hofmann et al. (2020); Bultinck et al. (2020); Kwan et al. (2021a); Khalaf et al. (2021); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Kwan et al. (2021b); Vafek and Kang (2021); Peri et al. (2021); Liu et al. (2021); Rossi (2021); Angeli and MacDonald (2021); Wang et al. (2022); Hu et al. (2022); Huhtinen et al. (2022); Törmä et al. (2022); Chan et al. (2022a); Herzog-Arbeitman et al. (2022a); Mera and Mitscherling (2022); Pan et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Lin et al. (2022); Herzog-Arbeitman et al. (2022, 2022); Peotta (2022); Khalaf et al. (2022); Chan et al. (2022b); Kang and Vafek (2019); Hofmann et al. (2023); Wan et al. (2023); Swaminathan et al. (2023); Hu et al. (2023); Löthman and Black-Schaffer (2017); Chen and Law (2023); Iskin (2023a); Peotta et al. (2023); Iskin (2023b); Tam and Peotta (2023); Zhang et al. (2023); Pan et al. (2023); Crépel et al. (2023); Mao et al. (2023); Kwan et al. (2023); Sahay et al. (2023); Jiang and Barlas (2023); Scheer and Lian (2023a, b); Devakul et al. (2023); Herzog-Arbeitman et al. (2023); Jia et al. (2023); Wu et al. (2023); Chen and Huang (2023); Salerno et al. (2023); Huang et al. (2022); Yu et al. (2023); Kashihara et al. (2023); Soldini et al. (2023); Jiang et al. (2023); Yu et al. (2023); Peralta Gavensky et al. (2023); Setty et al. (2023); Verma et al. (2023); Setty et al. (2023a).
@@
In this paper, we propose such a concept, which we call “Quantum Geometric Nesting” (QGN), that focuses on the geometric aspects of the electronic structure in flat-band systems. In contrast to Fermi surface nesting, here we start from a strong-coupling perspective in which the band dispersion is assumed negligible in comparison with the interaction strength, and then ask: for a given quantum geometry of the flat bands, are there any natural forms of order? If so, what sort of interactions favor them? We will show that these questions can be addressed by examining whether a set of QGN conditions is satisfied. Intuitively, these conditions refer to certain forms of ‘overlap’ of the wavefunctions in the flat band subspace at wavevectors related by a shift (p-h case) or inversion (p-p case) in momentum space, akin to the familiar case of Fermi surface nesting. For “perfectly nested quantum geometry,” we construct ideal Hamiltonians containing arbitrary combinations of an infinite class of engineered short-range two-body interactions that have solvable ground states with fermion bilinear order of a character determined by the quantum geometry. For these ideal Hamiltonians, the few-body excitations are solvable as well, including single-particle (quasi-particle), particle-hole (exciton), and two-particle (Cooper pair) excitations. For the p-p case, we further show that these ideal Hamiltonians host a (fine-tuned) pseudospin $SU(2)$ symmetry, whose stiffness can be inferred from the excitation spectra. We prove that certain symmetries (e.g. chiral or time-reversal) can ensure particular forms of QGN in flat bands. Non-symmetry-related cases also exist, such as an engineered model for pair-density-wave order that we construct. Extending the construction scheme, we further obtain an SSH-type model with solvable $p$-wave pairing groundstates.
Indeed, except for symmetry-enforced cases, it is unlikely for a physical system to have exact QGN - in common with the case of Fermi surface nesting. However, the use of QGN can be much more versatile: for systems with almost perfect QGN (quantification of the closeness to the ideal nesting will be defined below as ‘nestability’), one may engineer an idealized electronic structure with perfect QGN and then construct ideal interacting Hamiltonians that are close to the physical one. The solvability of these ideal models thus makes them excellent starting points for including perturbative corrections to reach the physical Hamiltonian. This type of intrinsically strong coupling viewpoint (in the sense that the dominant part of the interactions has been accounted for in the solvable models) has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al. (2016); Kwan et al. (2021b); Vafek and Kang (2021); Kang and Vafek (2019); Khalaf et al. (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Pan et al. (2022); Lin et al. (2022); Herzog-Arbeitman et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Crépel et al. (2023); Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023); Kwan et al. (2023). It is noteworthy that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems featuring fermion bi-linear order. We hence expect that QGN can greatly facilitate studies within this paradigm, in a way similar to the roles played by ‘ideal band conditions’ Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) in studies of fractional Chern insulator states.
@@
Figure 1: Schematic flowchart illustrating the structure of the theory: Given any electronic structure containing isolated flat bands, one can test whether a QGN condition is satisfied; if it is, one or more fermion bilinear order parameters (OP) will be suggested (Sec. I). For each OP, infinitely many “ideal” interacting Hamiltonians can be systematically constructed (Sec. II), which, after projecting onto the flat-band Hilbert space $\mathcal{H}_{\text{flat}}$, feature solvable ground states with the corresponding order (Sec. III) and solvable aspects of the collective excitation spectrum (Sec. IV). A practical recipe for using QGN in perturbative analyses of physical systems is proposed in Sec. II. Some examples of electronic structures with perfect QGN can be found in Sec. V. Possible extensions of this approach are commented on in Sec. VI.
@@ I.1 Preliminaries
We consider an electronic structure encoded in a tight-binding model  [^1] One could also formulate the model in continuum and work in momentum space throughout the discussion; no conclusions will be affected so long as lattice is not considered. Here we start with an explicit tight-binding model to avoid the subtleties when discussing the locality of the interactions and to connect to some existing literature. with charge conservation and translation symmetry among $\mathsf{V}$ unit cells in arbitrary dimension:
@@
where $\bm{R}$ indicates the unit cell position, Greek letters in lowercase $\mu,\nu,\dots$ denote a “flavor” index which, for simplicity, includes both orbital and spin indices unless otherwise specified. The annihilation operators for Bloch states with a given flavor index are expressed as $\hat{c}_{\bm{k},\mu}\equiv\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{R}}\hat{c}_{\bm{R},\mu}/\sqrt{\mathsf{V}}$ in the periodic embedding.  [^2] QGN can be defined for any embedding of the intra-unit-cell positions $\{\bm{r}_{\mu}\}$, but the periodic embedding $\{\bm{r}_{\mu}=\bm{0}\}$ is a special one since if a certain QGN condition is satisfied for any embedding, it is also satisfied in the periodic embedding. It thus suffices to always work with periodic embedding. In the sense that the embedding can be fixed, QGN is an embedding-independent concept, which is necessary for QGN to be a physical concept since embedding is purely a redundancy in the definition of Bloch basis Huhtinen et al. (2022). In momentum space, the single particle Hamiltonian is diagonalized with a band basis
@@
We will quantify the degree of QGN for flat-band systems by utilizing the geometric information in $U_{\mu n}(\bm{k})$. To do so, we consider an idealized setup (illustrated in Fig. 1) where there are $\mathsf{N}_{\text{flat}}$ nearly degenerate and flat bands isolated from all the other bands by a large gap $\Delta$, such that the relevant interaction strength $V$ is small compared to $\Delta$ but large compared to the flat bands’ bare bandwidth $W$,
@@ I.2 General Definitions
In terms of the gauge-invariant projection matrices $P$ and $Q$, the key information of QGN at a given wavevector $\bm{Q}$ for p-p or p-h channel is encoded in the linear operator
@@
where means complex conjugation. When viewed as a matrix (with $\mu\nu$ a grouped index and $\mu^{\prime}\nu^{\prime}$ the other), $\Pi$ is hermitian and positive semi-definite and can be diagonalized with non-negative eigenvalues.
By definition, an electronic structure has “perfect QGN” at a wavevector $\bm{Q}$ in the p-p or p-h channel when the corresponding $\Pi$ has eigenvalue(s) equal to 00. In other words, a perfect QGN is satisfied if there exists a $\mathsf{N}\times\mathsf{N}$ matrix $N^{\bm{Q}}_{\mu\nu}$, which we call the “nesting matrix”, satisfying (for any $\mu^{\prime}\nu^{\prime}$)
@@
where $F(\bm{k})$ is the form factor in the band basis corresponding to $N^{\bm{Q}}$ being the unprojected form factor in the original orbital basis,
@@
One important property (which can actually serve as an equivalent definition) of QGN is that for arbitrary $\bm{k}$, the form factor $F_{nm}(\bm{k})$ defined by a nesting matrix is a block-diagonal matrix that disconnects flat band indices from the others, i.e. $F_{nm}(\bm{k})=0$ as long as $n,m$ are not simultaneously $\leq\mathsf{N}_{\text{flat}}$ or $>\mathsf{N}_{\text{flat}}$. In other words, QGN implies that the wavefunctions of the flat bands at two momenta related by a shift (p-h case) or inversion (p-p case) are fully compatible, in the sense $N^{\bm{Q}}$ maps the single-particle flat band null spaces to each other. As a direct result, the static susceptibility of the OP defined in Eqs. 8&9 can be shown (see Sec. I C of SM) to strictly saturate an upper bound when the corresponding QGN is satisfied, suggesting the harmony between the quantum geometry and the OP in the ideal circumstances. We will further show in Sec. II that an infinite class of interacting solvable models can be constructed, which host ground states with the corresponding order.
For QGN to be non-trivial in the p-p channel, it is additionally required that $N^{\bm{Q}}$ cannot be symmetric since $\hat{O}^{\text{p-p},\bm{Q}}$ vanishes in this case. (It is proven by explicit construction that valid, anti-symmetric solutions can indeed exist for $\Pi$.) On the other hand, the identity matrix is always an allowed solution for Eq. 7 in the p-h case for $\bm{Q}=\bm{0}$, but this is a trivial one in the sense that the resulting $\hat{O}^{\text{p-h},\bm{Q}=\bm{0}}$ is simply the total density operator not associated with any broken symmetry. Excluding all the unwanted solutions, for a generic flat-band system, the value of the minimal eigenvalue of $\Pi^{\bm{Q}}$, $\varpi_{0}^{\bm{Q}}$, is a natural dimensionless measure of “nestability” that quantifies the ‘distance’ of a system to perfect QGN in the corresponding channel at $\bm{Q}$  [^4] We note that these concepts are spiritually similar to the “ideal band condition” Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) which are concerned with whether a Chern band is suitable for hosting fractional quantum Hall states.. Later in Sec. V we will give several examples of electronic structures with perfect QGN, i.e. $\varpi_{0}^{\bm{Q}}=0$.
@@
We mention that each satisfied QGN condition allows for constructing a generalized quantum geometry in the sense that it defines a notion of distance in momentum space and thus a generalized metric tensor, which can be related to the second spatial moment of the OP density bare correlation functions originating from the flat bands. The familiar definition of quantum distance $d(\bm{k},\bm{k}^{\prime})\equiv\sqrt{\mathsf{N}_{\text{flat}}-{\rm tr}{\left[P(\bm{k})P(\bm{k}^{\prime})\right]}}$ Provost and Vallee (1980); Resta (2011) using the Hilbert-Schmidt norm turns out to be one special case in this class of constructions, which is always valid since it corresponds to the ‘trivial QGN’ we mentioned above with $N^{\text{p-h},\bm{Q}=\bm{0}}=\mathbbm{1}$. We refer interested readers to Sec. I D of SM for formal definitions and discussions.
@@ II Ideal interacting Hamiltonians
For generic flat-band systems, no set of local hermitian operators that satisfy Eq. 14 exists. However, when the OP derives from a band geometry with perfect QGN, there is (at least) an infinite class of solutions that take a special flavor-space separable form
$\hat{S}^{(I)}_{\bm{R}}=$ $\sum_{\bm{R}_{1}\bm{R}_{2};\mu\nu}A^{(I)}(\bm{R}_{1},\bm{R}_{2})B^{(I)}_{\mu\nu}\hat{c}^{\dagger}_{\bm{R}+\bm{R}_{1},\mu}\hat{c}_{\bm{R}+\bm{R}_{2},\nu},$ (15)
where $A^{(I)}(\bm{R}_{1},\bm{R}_{2})=[A^{(I)}(\bm{R}_{2},\bm{R}_{1})]^{\star}$ is a spatial coefficient and $B^{(I)}$ is a hermitian matrix acting on orbital indices. In order for Eq. 14 to hold, the constraints on the coefficient $A^{(I)}$ and matrix $B^{(I)}$ are not overly restrictive, in that they only need to obey
$A^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=\left|A^{(I)}_{\bm{R}_{1},\bm{R}_{2}}\right|\mathrm{e}^{-\mathrm{i}(\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}/2},$ (16)
@@
for the p-h case. We derive in Sec. II B&C of SM a systematic way of constructing all possible $B^{(I)}$ and prove that there are at least $3\mathsf{N}_{\text{flat}}$ of them in the p-p case, and $\mathsf{N}_{\text{flat}}$ of them in the p-h case when $2\bm{Q}=\bm{0}$. (Some explicit worked examples are presented in Sec. V.) In particular, the identity matrix is always a solution for $B$ in the p-h case, which makes $\hat{S}_{\bm{R}}$ a local density operator for the unit cell at $\bm{R}$. Clearly, even given all the constraints, we are still left with an infinite class of possible interaction terms with considerable leeway, at least in choosing $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ and $A^{(I)}(\bm{R}_{1},\bm{R}_{2})$.
Although the above construction scheme could encompass many realistic forms of interactions, it is admittedly unlikely that the electronic structure and interactions of any physical system are perfectly ideal. Therefore, the merit of QGN should be viewed in a more practical way: it can help identify solvable limits of physical flat band systems of interest. Specifically, we are proposing the following recipe for the use of QGN in the study of a physical Hamiltonian $\hat{H}_{\text{phys}}$: First, one computes its electronic structure and examine whether there is a channel and a wavevector on which QGN is nearly satisfied (i.e. the corresponding $\varpi_{0}$ is small compared to $1$). If there is indeed such a nearly nested possibility, one can consider its perfect nesting limit by slightly modifying the electronic structure either numerically or analytically (through recognizing certain approximate symmetries, e.g. those discussed below in Sec. V). Then, one can further search in the space of all possible ideal solvable interacting Hamiltonians validated by the QGN, to find one that is closest to the physical Hamiltonian, $\hat{H}_{\text{ideal}}$. This can be systematically achieved by minimizing the norm of the difference between them, $||\delta\hat{H}||=||\hat{H}_{\text{ideal}}-\hat{H}_{\text{phys}}||$. A small value of this operator distance justifies perturbative analyses starting from the ideal solvable limit. The flexibility of our construction allows the dominant part of many interactions to be captured. Furthermore, we remark that these nearby solvable limits can also facilitate numerical studies of $\hat{H}_{\text{phys}}$, in that one can benchmark the validity of the methods being used at these limits, and confirm the phase being observed by checking whether there is an intervening phase transition as interpolating between $\hat{H}_{\text{phys}}$ and $\hat{H}_{\text{ideal}}$.
One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”. For the special case of Hubbard interactions, a property called the “uniform pairing condition” Tovmasyan et al. (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al. (2022). It is possible to generalize this condition, which we will exemplify in Sec. V.3. However, these quadratic terms are not of central focus since the current work is concerned with the existence of ideal interacting Hamiltonians for a given set of isolated flat bands, whose quantum geometry is encoded in the projection matrix and is unaffected by these terms to the leading order in $V/\Delta$. An operational workaround is to simply regard these additional quadratic terms as parts of $\hat{H}_{0}$. For more detailed discussions about this nomenclature subtlety, see Sec. II D of SM.
We remark that the building blocks of the ideal Hamiltonians, $\{\hat{S}^{I}_{\bm{R}}\}$, do not necessarily mutually commute. As we will see in the next section, the solvability of these interactions results from the fact that they all commute with the order parameter and share certain eigenstates. The ideal interacting Hamiltonians are in this sense ‘frustration-free’.